Prove that the given equations are identities:
a)cos (30°+x)+ cos (150°-x)=0
b)cos (60°+x)- cos (300°-x)=0

Answer by jsmallt9(3758)

Both of these problems require the sum and difference formulas for cos:
cos(a+b) = cos(a)*cos(b)-sin(a)*sin*(b)
cos(a-b) = cos(a)*cos(b)+sin(a)*sin*(b)

a) cos (30°+x)+ cos (150°-x)=0

(cos(30)*cos(x) - sin(30)*sin(x)) + (cos(150)*cos(x) + sin(150)*sin(x)) = 0

And since everything cancels out...
0+=+0

b)cos (60°+x)- cos (300°-x)=0
(cos(60)cos(x) - sin(60)*sin(x)) - (cos(300)*cos(x) + sin(300)sin(x)) = 0

Note the "-" in front of the second "half" of the left side of the equation. As a result, everything cancels out (again) leaving...
0+=+0